# A Multimodal Dataset for Visually Grounded Ambiguity in Machine Translation

### 03 May 2026 v1

@@ Abstract
Ambiguity resolution is a key challenge in multimodal machine translation (MMT), where models must genuinely leverage visual input to map an ambiguous expression to its intended meaning. Although prior work has proposed disambiguation-oriented benchmarks that provide supportive evidence for the role of vision, we observe substantial issues in data quality and a mismatch with translation scenarios. Moreover, existing ambiguity-oriented evaluations are not well suited to broader ambiguity types in open-ended translation. To address these limitations, we present VIDA (Visually-Dependent Ambiguity), a dataset of 2,500 carefully curated instances in which resolving an annotated ambiguous source span requires visual evidence. We further propose Disambiguation-Centric Metrics that use an LLM-as-a-judge classifier to verify whether annotated ambiguous expressions are resolved correctly at the span level. Experiments with two state-of-the-art Large Vision Language Models under vanilla inference, supervised fine-tuning (SFT), and our chain-of-thought SFT (CoT-SFT) show that while SFT improves overall translation quality, CoT-SFT yields more consistent gains in disambiguation accuracy, especially on out-of-distribution subsets, indicating a stronger generalization for resolving diverse ambiguity types.
@@ 1 Introduction
Multimodal machine translation (MMT) extends neural machine translation by incorporating visual context to improve translation quality (Lala and Specia, 2018; Yao and Wan, 2020). Recent Large Vision Language Models (LVLMs) show impressive performance on MMT benchmarks (Bai et al., 2025; Zhu et al., 2025). However, it is still unclear whether LVLMs truly leverage visual information during translation. For instance, prior studies (Elliott, 2018; Wu et al., 2021) show that replacing or perturbing images often leads to only minor degradations, raising questions about the actual contribution of the visual modality.
@@
To investigate this question, recent works (Ma et al., 2024; Wang et al., 2024a) have adopted ambiguity resolution as a probe of visual dependence in MMT, introducing benchmarks such as 3AM (Ma et al., 2024) and MMA (Wang et al., 2024a). 3AM targets English–Chinese translation and primarily focuses on word-level ambiguity, but contains many instances that remain resolvable from text alone. MMA focuses on sentence-level ambiguity in a VQA-style format, which is not directly aligned with the translation scenario. Additionally, prior work on MMT ambiguity evaluation spans several paradigms, including contrastive evaluation with predefined translation candidates (Futeral et al., 2023), as well as task-specific metrics based on predefined variants or rule-based matching for narrower ambiguity phenomena (Lala and Specia, 2018; Li et al., 2021). While these settings provide ambiguity-oriented evaluation, they are less suitable for open-ended MMT generation with broader ambiguity types and valid paraphrases. General MT metrics such as BLEU and COMET are also useful for overall translation quality, but they are not designed to measure whether a particular ambiguous span has been resolved correctly.
@@
To reliably evaluate visual disambiguation in MMT, we need both an MMT dataset with visually-dependent ambiguities and an evaluation method that directly measures disambiguation accuracy. In this work, we address these limitations by introducing VIDA (Visually-Dependent Ambiguity), a dataset of 2,500 instances curated to capture visually-dependent translation ambiguities. VIDA comprises three subsets: VIDA-Base, VIDA-Sent, and VIDA-CollN, covering word- and sentence-level ambiguities, as well as collective-noun cases where visual context is necessary to translate abstract entities into concrete target expressions. We further propose Disambiguation-Centric Metrics based on an LLM-as-a-judge classifier. Unlike standard translation metrics, our metrics explicitly target disambiguation accuracy and provide a more sensitive measure of correct ambiguity resolution.
@@
Beyond dataset and evaluation, we further explore a chain-of-thought supervised fine-tuning (CoT-SFT) strategy (Muennighoff et al., 2025) to examine whether reasoning supervision can elicit visually grounded disambiguation in MMT. Specifically, we manually synthesize task-specific reasoning traces that guide the model to identify the ambiguous expression and justify its resolution with visual evidence before generating the final translation. Experiments show that CoT-SFT yields consistent gains in disambiguation accuracy over standard SFT, suggesting better generalization across diverse ambiguity types. The main contributions of this paper are as follows.
@@ 2 Related Works
Recent benchmarks such as 3AM (Ma et al., 2024) and MMA (Wang et al., 2024a) address multimodal disambiguation but show limitations in MMT evaluation. 3AM contains instances that are resolvable by text alone, reducing visual sensitivity to visual grounding, while MMA uses a VQA-style format incompatible with translation scenarios. These limitations motivate our VIDA dataset, which covers both word- and sentence-level cases strictly requiring visual evidence for correct interpretation.
Evaluation Metrics for MMT.
Prior work on multimodal ambiguity evaluation has used several different settings. CoMMuTE (Futeral et al., 2023) evaluates disambiguation with predefined translation candidates rather than open-ended translation generation, while MLT (Lala and Specia, 2018) and AmbigCaps (Li et al., 2021) evaluation focus on word-level ambiguity and rely on rule-based string matching. Furthermore, standard MT metrics such as BLEU (Papineni et al., 2002) and COMET (Rei et al., 2020) do not directly verify whether an ambiguous span has been resolved correctly, since surface-overlap metrics may penalize valid paraphrases or lexical variation and sentence-level metrics are too coarse-grained for span-level disambiguation. In this work, we propose Disambiguation-Centric Metrics using an LLM-as-a-judge classifier to directly assess span-level disambiguation accuracy in open-ended MMT generation.
@@ 3 Dataset Curation
To construct MMT dataset with visually-dependent translation ambiguities, we adopt a three-stage semi-automatic pipeline, as shown in Figure 1. Our goal is to collect instances where the disambiguating evidence is grounded in the image, and to provide human-verified reference translations.
@@ 3.1 Stage 1: Data Preprocessing and Filtering
The goal of this stage is to curate image–text aligned and textually clean source captions with visually-dependent ambiguities. We use GPT-4o (Hurst et al., 2024) to filter mismatched image–text pairs and normalize text. We then apply a dual-model consensus strategy with two state-of-the-art commercial LLMs, Qwen-Max (Team, 2024) and DeepSeek-v3 (Liu et al., 2024): each model independently assesses whether the caption remains ambiguous without the image. We retain only captions that both models judge as ambiguous, and record an Ambiguous Caption and an Ambiguity Rationale that points the ambiguous span and the disambiguation cue.
@@ 3.3 Stage 3: Quality Assurance and Validation
To ensure translation correctness, we further conduct human verification with two native Chinese speakers (one Ph.D. and one M.S. in Computer Science). For each instance, annotators are shown the image, the source caption, the ambiguity rationale, the candidate translation, and Qwen-Max’s judgment. Each annotator independently reviews the translation using three criteria: (i) whether the annotated ambiguity is correctly resolved, (ii) whether the translation is fluent, and (iii) whether the meaning of the translation is preserved. If the candidate fails any criterion, the annotator provides a corrected translation based on the image and the ambiguity rationale. During verification, we identified collective noun cases where underspecified terms must be concretized based on the image content and group these cases into the Collective Noun Subset. Finally, VIDA comprises 2,500 instances across three subsets. See Appendix A for details.
@@ 4 Evaluation Metrics for Disambiguation
In MMT disambiguation, the objective of evaluation is to measure disambiguation accuracy rather than overall translation quality. We use an LLM-as-a-judge to classify whether the ambiguous span is correctly resolved, and compute disambiguation accuracy from the binary outputs. Specifically, we fine-tune Qwen3-8B on VIDA to determine whether annotated ambiguous expressions are correctly resolved in translation. We train the classifier in a contrastive setting. Gold disambiguated translations are used as positive samples. Negative samples are candidate translations from our curation pipeline that fails to resolve the annotated ambiguity. Building on this, we introduce two complementary Disambiguation-Centric Metrics for a comprehensive evaluation of disambiguation performance: Disambi-Term measures term-level disambiguation accuracy by evaluating each annotated ambiguous term in the entire dataset. Disambi-Inst. reports instance-level accuracy, counting a sentence as correct only if all ambiguous expressions within it are correctly resolved.
@@ 5.1 Experimental Settings
We conduct experiments on the VIDA dataset. VIDA-Base is split into training and test sets with a 7:3 ratio, resulting in 1,352 training samples and 580 test samples, Due to their limited size, VIDA-CollN and VIDA-Sent are used exclusively as out-of-distribution (OOD) test sets.
@@
We report standard translation metrics to assess overall translation quality, including BLEU, chrF (Popović, 2015), chrF++ (Popović, 2017), METEOR (Banerjee and Lavie, 2005), TER (Snover et al., 2006), BERT-F1 (Devlin et al., 2019) and COMET. More importantly, we evaluate disambiguation performance using the proposed Disambiguation-Centric Metrics, Disambi-Term and Disambi-Inst.
@@
We evaluate two state-of-the-art LVLMs: Qwen2.5-VL-7B (Bai et al., 2025), InternVL3-8B (Zhu et al., 2025). For each model, we consider three settings: (i) the vanilla model without task-specific fine-tuning, (ii) supervised fine-tuning (SFT), and (iii) the Chain-of-Thought Supervised Fine-Tuning (CoT-SFT), which augments SFT with manually synthesized reasoning traces for ambiguity resolution (Appendix C).
@@ 5.2 Experimental Results
Table 1 summarizes results for two LVLMs under Vanilla, SFT, and CoT-SFT on in- and out-of-distribution tests.
Analysis on In-Distribution Dataset.
Evaluation on the in-distribution dataset (VIDA-Base-Test) serves to examine how well the models adapt to the training distribution (VIDA-Base-Train). On VIDA-Base-Test, SFT setting achieves the strongest overall translation quality for both models across most standard metrics. CoT-SFT remains competitive on semantic metrics (e.g., COMET) for both models, but can slightly underperform SFT on surface-overlap metrics such as BLEU, likely because ambiguity resolution often requires paraphrasing or structural reformulation that deviates from reference wording. Under Disambiguation-Centric evaluation, CoT-SFT yields the best performance for InternVL3-8B (Disambi-Term/Inst. 64.89/51.38), surpassing both SFT and Vanilla. In contrast, for Qwen2.5-VL-7B, CoT-SFT improves over Vanilla but remains slightly below SFT (60.71/46.90 vs. 61.42/49.31), suggesting that in-distribution gains from reasoning supervision are model-dependent. We attribute part of the CoT-SFT disambiguation gap to overthinking behaviors and provide qualitative analysis in Appendix  D.
@@
Evaluation on the out-of-distribution (OOD) datasets (VIDA-Sent and VIDA-CollN) tests whether models can generalize the disambiguation to unseen ambiguity types. Compared to in-distribution results, CoT-SFT shows clearer advantages under distribution shift. For standard translation metrics, CoT-SFT achieves the highest COMET on both OOD subsets for both models, indicating stronger semantic adequacy on unseen ambiguity types. Disambiguation-centric metrics make the advantage more explicit. Compared to SFT, CoT-SFT achieves consistent gains on VIDA-Sent (approximately +3.5 points for InternVL3-8B and +7.5 points for Qwen2.5-VL-7B on both Disambi-Term and Disambi-Inst.), and the improvements are substantially larger on VIDA-CollN (over +15 points for InternVL3-8B and over +12 points for Qwen2.5-VL-7B), suggesting that CoT-SFT demonstrates stronger generalization beyond the training distribution.
@@
All-Test merges all subsets into a single evaluation rather than averaging their scores, reflecting overall performance weighted by subset sizes. On standard translation metrics, SFT consistently improves over Vanilla for both models, indicating better overall translation quality on the mixed test distribution. In contrast, CoT-SFT consistently yields higher semantic adequacy and disambiguation than SFT: it achieves the highest COMET and improves Disambi-Term/Inst. by about +4.1/+5.0 for InternVL3-8B and +2.7/+3.6 for Qwen2.5-VL-7B. Overall, CoT-SFT achieves a strong balance between semantic adequacy and disambiguation and transfers better to the diverse ambiguity types.
@@ 6 Conclusion
We introduce VIDA, an MMT dataset with visually-dependent translation ambiguities covering both word- and sentence-level, and propose Disambiguation-Centric Metrics that directly measure disambiguation accuracy using an LLM-as-a-judge classifier. We further compare CoT-SFT against SFT for disambiguation. Our findings show that SFT achieves strong overall translation quality on in-distribution data, whereas CoT-SFT achieves larger gains in disambiguation accuracy on the mixed All-Test evaluation, suggesting better generalization for resolving diverse ambiguity types.
@@
Although VIDA targets highly visually dependent ambiguities, the dataset remains modest in scale compared to large-scale translation corpora and is currently limited to English–Chinese translation. In addition, the reasoning traces used to supervise CoT-SFT are synthetically constructed based on manually designed patterns, which may not reflect natural model reasoning. Future work includes expanding the dataset to additional language pairs and exploring training paradigms that encourage models to develop task-relevant reasoning behaviors without relying on handcrafted traces, such as reinforcement learning or preference-based optimization methods.
@@
The proposed VIDA dataset is curated from publicly available data sources. The dataset will be released with usage guidelines to support research on multimodal machine translation and ambiguity resolution.
Appendix A Dataset Statistics
A.1 Pipeline Filtering Statistics
The large reduction from the raw candidate pool to the final benchmark reflects the strict quality control in our curation pipeline. We begin with 26,452 raw image–text pairs, retain 14,993 after image–text matching and text normalization, and then keep only 2,500 instances after ambiguity detection and dual-LLM verification. Subsequent stages refine translations and annotations without discarding additional instances. All 2,500 retained instances are then human-verified by two native Chinese annotators, corresponding to 100% human verification coverage. The filtering funnel is summarized in Table 2.
A.2 VIDA Subset Statistics
The rigorous pipeline outlined in section 3 results in the construction of a new dataset, VIDA (Visually-Dependent Ambiguity). In total, VIDA comprises 2,500 instances specifically curated to feature high ambiguity complexity and visual dependency. The dataset comprehensively covers both word-level and sentence-level ambiguities and is organized into the following three subsets:
  • VIDA-Base: Curated from the 3AM dataset, this subset contains 1,932 samples, primarily focusing on word-level ambiguities that require visual context for resolution.
  • VIDA-CollN (Collective Noun Subset): This specialized subset consists of 256 samples focusing on the disambiguation of collective nouns, where the abstract nature of the group is made concrete by the associated visual information.
  • VIDA-Sent: Adapted from the MMA dataset, this subset provides 312 samples. These instances tend to exhibit more complex, sentence-level semantic ambiguities that necessitate a holistic understanding of the image for correct interpretation and translation.
A complete statistical summary of these subsets is provided in Table 3. For each subset, the column Ambiguity specifies the primary ambiguity focus (word-level, sentence-level, or mixed). Size denotes the total number of samples in each subset. Avg. Length (Words) gives the average sentence length measured by the number of words, and Avg. Ambi. indicates the average number of ambiguities per sentence.
VIDA-Base is the largest subset (1,932 samples), consisting primarily of word-level ambiguities. It contains relatively longer sentences, averaging 11.12 words, and exhibits the highest ambiguity density (1.78 ambiguous terms per sentence). VIDA-Sent (312 samples) specifically focuses on sentence-level ambiguities, with shorter sentences averaging 6.00 words and exactly one annotated ambiguity per instance. Finally, VIDA-CollN (256 samples) also targets word-level ambiguities, specializing in collective nouns. Compared to VIDA-Base, VIDA-CollN features shorter sentences (10.08 words on average) and a lower ambiguity density (1.20 per sentence).

### 26 May 2026 v2

# VIDA: A Dataset for Visually Dependent Ambiguity in Multimodal Machine Translation

@@ Abstract
Ambiguity resolution is a key challenge in multimodal machine translation (MMT), where models must genuinely leverage visual input to map an ambiguous expression to its intended meaning. Although prior work has proposed disambiguation-oriented benchmarks probing the role of vision, we observe that existing benchmarks remain limited by task-format mismatch, narrow ambiguity coverage, or insufficient visual-dependency validation. Moreover, existing ambiguity evaluations are not well suited to diverse ambiguity types in open-ended translation. To address these limitations, we present VIDA (Visually-Dependent Ambiguity), a dataset of 2,500 carefully curated instances in which resolving an annotated source span requires visual evidence. We further propose Disambiguation-Centric Metrics that use an LLM-as-a-judge classifier to verify whether annotated ambiguous expressions are resolved correctly at the span level. Experiments with two state-of-the-art LVLMs show that supervised fine-tuning (SFT) improves overall translation quality, while chain-of-thought SFT (CoT-SFT) yields stronger out-of-distribution disambiguation, suggesting that explicit disambiguation guidance improves generalization to diverse ambiguity types.
@@ 1 Introduction
Multimodal machine translation (MMT) extends neural machine translation by incorporating visual context to improve translation quality (Lala and Specia, 2018; Yao and Wan, 2020). Recent Large Vision Language Models (LVLMs) show impressive performance on MMT benchmarks (Bai et al., 2025; Zhu et al., 2025). However, it remains unclear whether LVLMs truly leverage visual information during MMT. Earlier studies (Elliott, 2018; Wu et al., 2021) showed that replacing or perturbing images often leads to only minor degradations, raising questions about the actual contribution of the visual modality. This concern remains relevant in the LVLM era: stronger visual perception does not necessarily guarantee reliable use of visual evidence during translation.
@@
Ambiguity resolution provides a direct probe of visual dependence in MMT, and existing work has made progress in this direction by introducing multimodal disambiguation benchmarks (Ma et al., 2024; Wang et al., 2024a; Futeral et al., 2023). Yet no existing benchmark combines open-ended MMT generation with broad ambiguity coverage and visual-dependency validation. CoMMuTE (Futeral et al., 2023) is a contrastive benchmark for word-level and gender ambiguity that scores candidate translations rather than requiring models to generate translations freely. 3AM (Ma et al., 2024) targets English–Chinese MMT and primarily focuses on word-level ambiguity, but includes many instances resolvable from text alone. MMA (Wang et al., 2024a) covers sentence-level ambiguity in a VQA-style format rather than translation generation. Additionally, prior work on MMT ambiguity evaluation spans several paradigms, including contrastive evaluation with predefined translation candidates (Futeral et al., 2023) and rule-based matching with lexical variants (Lala and Specia, 2018; Li et al., 2021). While these settings provide ambiguity-oriented evaluation, they are less suitable for open-ended MMT generation with broader ambiguity types and valid paraphrases. General MT metrics such as BLEU and COMET assess overall translation quality, but do not directly indicate whether a particular ambiguous span has been resolved correctly.
@@
Reliable evaluation of visual disambiguation in MMT requires both an MMT dataset with visually-dependent ambiguities and an evaluation method that directly measures disambiguation accuracy. We therefore introduce VIDA (Visually-Dependent Ambiguity), an MMT dataset of 2,500 instances capturing visually-dependent translation ambiguities at the word and sentence levels, as well as collective-noun phenomena. We further propose Disambiguation-Centric Metrics based on an LLM-as-a-judge classifier. These metrics complement standard translation metrics by explicitly targeting span-level disambiguation accuracy.
@@
Beyond dataset and evaluation, we use chain-of-thought supervised fine-tuning (CoT-SFT) (Muennighoff et al., 2025) with a manually designed template as a diagnostic setting to test whether explicit visual reasoning supervision improves disambiguation beyond standard SFT. Experiments show that CoT-SFT yields stronger disambiguation under OOD and aggregate evaluation, suggesting better generalization across diverse ambiguity types. The main contributions of this paper are:
@@ 2 Related Work
Prior work has studied visual disambiguation through translation-based benchmarks (Lala and Specia, 2018; Li et al., 2021; Ma et al., 2024), contrastive MMT benchmarks (Futeral et al., 2023), and VQA-style ambiguity benchmarks (Wang et al., 2024a). In comparison, VIDA targets open-ended MMT generation with broader ambiguity coverage, retaining only cases whose correct interpretation requires visual evidence.
Disambiguation Metrics for MMT.
Prior work has evaluated multimodal ambiguity with contrastive candidate ranking (Futeral et al., 2023), rule-based matching over predefined lexical variants (Lala and Specia, 2018; Li et al., 2021), and general MT metrics (Papineni et al., 2002; Rei et al., 2020). These approaches are less suitable for open-ended MMT generation with broader ambiguity types, where correct disambiguation may be expressed through paraphrases or lexical variation. We therefore propose Disambiguation-Centric Metrics using an LLM-as-a-judge classifier to directly assess span-level disambiguation accuracy. A detailed comparison is provided in Appendix A.
@@ 3 Dataset Curation
To construct an MMT dataset with visually-dependent translation ambiguities, we adopt a three-stage semi-automatic pipeline, as shown in Figure 1. Our goal is to collect instances where the disambiguating evidence is grounded in the image, and to provide human-verified reference translations.
@@ 3.1 Stage 1: Data Preprocessing and Filtering
The goal of this stage is to curate image–text aligned and textually clean source captions with visually-dependent ambiguities. We use GPT-4o (Hurst et al., 2024) to filter mismatched image–text pairs and normalize text. We then apply a dual-model consensus strategy with two state-of-the-art commercial LLMs, Qwen-Max (Team, 2024) and DeepSeek-v3 (Liu et al., 2024): each model independently assesses whether the caption remains ambiguous without the image. We retain only captions that both models judge as ambiguous, and record an Ambiguous Caption and an Ambiguity Rationale that points to the ambiguous span and the disambiguation cue.
@@ 3.3 Stage 3: Quality Assurance and Validation
To ensure translation correctness, we further conduct human verification with two native Chinese speakers (one Ph.D. and one M.S. in Computer Science). For each instance, annotators are shown the image, the source caption, the ambiguity rationale, the candidate translation, and Qwen-Max’s judgment. Each annotator independently reviews the translation using three criteria: (i) whether the annotated ambiguity is correctly resolved, (ii) whether the translation is fluent, and (iii) whether the meaning of the translation is preserved. If the candidate fails any criterion, the annotator provides a corrected translation based on the image and the ambiguity rationale. During verification, we identified collective noun cases where underspecified terms must be concretized based on the image content and group these cases into the Collective Noun Subset. Finally, VIDA comprises 2,500 instances across three subsets (VIDA-Base, VIDA-Sent, and VIDA-CollN). See Appendix C for details.
@@ 4 Evaluation Metrics for Disambiguation
In MMT disambiguation, the objective of evaluation is to measure disambiguation accuracy rather than overall translation quality. We use an LLM-as-a-judge to classify span-level correctness and compute accuracy from binary outputs. Specifically, we fine-tune Qwen3-8B on VIDA to determine whether annotated ambiguous expressions are correctly resolved in translation. We train the classifier in a contrastive setting. Gold disambiguated translations are used as positive samples. Negative samples are candidate translations from our curation pipeline that fail to resolve the annotated ambiguity. Building on this, we introduce two complementary Disambiguation-Centric Metrics for a comprehensive evaluation of disambiguation performance: Disambi-Term measures term-level disambiguation accuracy by evaluating each annotated ambiguous term in the entire dataset. Disambi-Inst. reports instance-level accuracy, counting a sentence as correct only if all ambiguous expressions within it are correctly resolved. Judge reliability and training details are provided in Appendix D.
@@ 5.1 Experimental Settings
We conduct experiments on the VIDA dataset. VIDA-Base is split 7:3 into 1,352 training and 580 test samples, while VIDA-CollN and VIDA-Sent are held out as out-of-distribution (OOD) test sets due to limited size.
@@
We report standard MT metrics (BLEU, chrF/chrF++ (Popović, 2015, 2017), METEOR (Banerjee and Lavie, 2005), TER (Snover et al., 2006), BERT-F1 (Devlin et al., 2019), and COMET) together with Disambi-Term and Disambi-Inst to evaluate disambiguation accuracy.
@@
We evaluate two state-of-the-art LVLMs: Qwen2.5-VL-7B (Bai et al., 2025), InternVL3-8B (Zhu et al., 2025), under three settings: vanilla inference, supervised fine-tuning (SFT), and chain-of-thought supervised fine-tuning (CoT-SFT) with manually synthesized ambiguity-resolution traces (Appendix E).
@@ 5.2 Experimental Results
Analysis on In-Domain Dataset.
Evaluation on the in-domain test set (VIDA-Base-Test) in Table 1 examines how well the models fit the training distribution (VIDA-Base-Train). The SFT setting achieves the strongest overall translation quality for both models across most standard metrics. CoT-SFT remains competitive on semantic metrics (e.g., COMET) for both models, but can slightly underperform SFT on surface-overlap metrics such as BLEU, likely because ambiguity resolution often requires paraphrasing or structural reformulation that deviates from reference wording. To further assess whether these surface-metric gaps indicate an overall quality loss, we use a commercial LVLM judge and find that CoT-SFT remains comparable to SFT in overall quality (Appendix B). Under Disambiguation-Centric evaluation, CoT-SFT yields the best performance for InternVL3-8B (Disambi-Term/Inst. 64.89/51.38), surpassing both SFT and Vanilla. In contrast, for Qwen2.5-VL-7B, CoT-SFT improves over Vanilla but remains slightly below SFT (60.71/46.90 vs. 61.42/49.31), suggesting that in-domain gains from reasoning supervision are model-dependent. We attribute part of the CoT-SFT disambiguation gap to overthinking behaviors and provide qualitative analysis in Appendix F.
@@
Evaluation on the out-of-distribution (OOD) datasets (VIDA-Sent and VIDA-CollN) examines whether models can generalize to unseen ambiguity types. Compared to in-domain results, CoT-SFT shows clearer advantages under distribution shift. For standard translation metrics, CoT-SFT achieves the highest COMET on both OOD subsets for both models, indicating stronger semantic adequacy on unseen ambiguity types. Disambiguation-centric metrics make the advantage more explicit. On VIDA-Sent, CoT-SFT achieves consistent gains over SFT (approximately +3.5 points for InternVL3-8B and +7.5 points for Qwen2.5-VL-7B on both Disambi-Term and Disambi-Inst.). The improvements are substantially larger on VIDA-CollN (over +15 points for InternVL3-8B and over +12 points for Qwen2.5-VL-7B), suggesting that CoT-SFT demonstrates stronger generalization beyond the training distribution.
@@
All-Test merges all subsets into a single evaluation rather than averaging their scores, reflecting overall performance weighted by subset sizes. On standard translation metrics, SFT generally improves over Vanilla for both models, indicating better overall translation quality on the mixed test distribution. In contrast, CoT-SFT consistently yields higher semantic adequacy and disambiguation accuracy than SFT: it achieves the highest COMET and improves Disambi-Term/Inst. by about +4.1/+5.0 for InternVL3-8B and +2.7/+3.6 for Qwen2.5-VL-7B. Overall, CoT-SFT balances semantic adequacy and disambiguation accuracy while generalizing better to diverse ambiguity types.
@@ 6 Conclusion
We introduce VIDA, an MMT dataset for visually-dependent word- and sentence-level ambiguities, and propose Disambiguation-Centric Metrics that directly measure disambiguation accuracy based on an LLM-as-a-judge classifier. Our comparison of SFT and CoT-SFT shows that SFT better fits the surface patterns of the training references, whereas CoT-SFT yields stronger visually grounded disambiguation, especially under distribution shift. This suggests that explicit disambiguation guidance improves generalization to diverse ambiguity types.
@@
Although VIDA targets highly visually dependent ambiguities, the dataset remains modest in scale compared to large-scale translation corpora and is currently limited to English–Chinese translation. In addition, the reasoning traces used to supervise CoT-SFT are synthetically constructed based on manually designed patterns, which may not reflect natural model reasoning. Our CoT-SFT experiments are also conducted on two mid-sized open-source LVLMs, which leaves open the question of how the strategy transfers to larger or proprietary LVLMs. Future work includes expanding the dataset to additional language pairs, validating the CoT-SFT findings on a broader range of LVLMs, and moving beyond fixed templates toward dynamic or adaptive reasoning generation, for instance through reinforcement learning or preference-based optimization methods that let the model adapt its reasoning depth to the difficulty of the ambiguity.
@@
The proposed VIDA dataset is curated from publicly available data sources and inherits their research-use license; the images do not introduce additional personally identifying information beyond the underlying datasets. Annotators were informed of the task goals before annotation. The dataset will be released with usage guidelines to support research on multimodal machine translation and ambiguity resolution.
Appendix A Positioning Against Prior Resources
Resource Generation format Ambiguity focus Ambiguity evaluation
MLT (Lala and Specia, 2018) localized lexical translation lexical choice rule-based lexical matching
AmbigCaps / M3-AmbigCaps (Li et al., 2021; Guo et al., 2022) open-ended caption translation primarily gender/pronoun ambiguity rule-based matching over lexical variants
CoMMuTE (Futeral et al., 2023) contrastive candidate ranking lexical/gender ambiguity contrastive accuracy over predefined candidates
3AM (Ma et al., 2024) open-ended MMT generation mainly word-level ambiguity general MT and ambiguity-oriented evaluation
MMA (Wang et al., 2024a) VQA-style disambiguation sentence-level ambiguity VQA-style answer accuracy
VIDA open-ended MMT generation lexical, sentence, pragmatic, collective noun LLM-judge semantic span accuracy
Table 2: Positioning of VIDA relative to prior multimodal disambiguation resources.
Existing resources address visual disambiguation from different angles, but they differ substantially from VIDA in task format and validation target. MLT focuses on localized lexical translation, where the model resolves an ambiguous source word with sentence and image context rather than generating a full translation. AmbigCaps and M3-AmbigCaps use open-ended caption translation, but their ambiguity coverage is concentrated on gender or pronoun choices derived from back-translation. CoMMuTE provides a controlled contrastive setting by asking models to rank predefined translation candidates, which is useful for isolating visual cues but does not evaluate free-form generation. 3AM targets English–Chinese MMT but primarily focuses on word-level ambiguity and includes many instances resolvable from text alone. MMA broadens the phenomenon to sentence-level ambiguity, but is formulated as a VQA task rather than translation. In contrast, VIDA targets open-ended MMT generation, covers word-, sentence-, pragmatic-, and collective-noun ambiguity, and retains only instances whose intended interpretation is verified to require visual evidence.
The evaluation protocols also differ in what they can verify. Contrastive accuracy measures whether a model ranks a predefined correct candidate above distractors, while rule-based matching over lexical variants is tied to expected surface forms. These designs are less suitable for open-ended translation, where correct disambiguation may be expressed through paraphrases or different lexical choices. General MT metrics such as BLEU and COMET remain useful for overall translation quality, but they do not directly verify whether an annotated ambiguous span is resolved correctly, since surface-overlap metrics may penalize valid paraphrases or lexical variation and sentence-level metrics are too coarse-grained for span-level disambiguation. VIDA therefore uses Disambiguation-Centric Metrics to judge span-level semantic resolution while allowing valid paraphrases. Table 2 summarizes the comparison.
Appendix B Auxiliary Overall Quality Evaluation
To complement standard MT metrics, we evaluate overall translation quality with a commercial LVLM judge on a 0–100 scale. Table 3 shows that SFT and CoT-SFT are close in overall quality, and CoT-SFT is often comparable or higher despite stronger disambiguation performance. This supports our main interpretation that CoT-SFT’s gains reflect improved ambiguity resolution rather than a severe loss in general translation quality.
Model Dataset Vanilla SFT CoT-SFT
InternVL3-8B All-Test 83.82 84.76 86.35
VIDA-Base-Test 86.06 87.24 87.25
VIDA-Sent 79.31 78.86 83.57
VIDA-CollN 84.26 86.34 87.70
Qwen2.5-VL-7B All-Test 83.92 85.70 85.87
VIDA-Base-Test 86.07 87.33 87.33
VIDA-Sent 79.10 82.24 82.48
VIDA-CollN 84.91 86.24 86.69
Table 3: Auxiliary overall translation quality judged by a commercial LVLM. Scores are on a 0–100 scale and supplement the main metrics in Table 1.
Appendix C Dataset Statistics